### v1

# A Cramér–Wold theorem for mixtures

Abstract.
We show how a Cramér–Wold theorem for a family of multivariate probability distributions can be used to generate a similar theorem for mixtures (convex combinations) of distributions drawn from the same family.
Using this abstract result, we establish a Cramér–Wold theorem for mixtures of multivariate Gaussian distributions. According to this theorem, two such mixtures can be distinguished by projecting them onto a certain predetermined finite set of lines, the number of lines depending only on the total number Gaussian distributions involved and on the ambient dimension. A similar result is also obtained for mixtures of multivariate $t$-distributions.
Key words and phrases:
Projection, identifiability, mixture, Gaussian distribution, $t$-distribution
2020 Mathematics Subject Classification:
60B11
Fraiman and Moreno were supported by grant FCE-3-2022-1-172289, Agencia Nacional de Investigación e Innovación, Uruguay. Ransford was supported by NSERC Discovery Grant RGPIN–2020–04263.
1. Introduction
The Cramér–Wold device is the name given to a general technique for analyzing multivariate probability distributions via their lower-dimensional projections. The name originates from a classical theorem of Cramér and Wold [3] to the effect that a probability measure in Euclidean $d$-dimensional space is uniquely determined by its one-dimensional projections in all directions.
For certain classes of measures, one can do better. Indeed, in some cases, just finitely many projections suffice to determine the measure. As an example, we cite the case of elliptic measures in $\mathbb{R}^{d}$ (which includes that of Gaussian measures), where just $(d^{2}+d)/2$ suitably chosen projections suffice (see [5]).
The central object of study in this article is that of mixtures, namely convex combinations of probability measures. For example, suppose that we already have a Cramér–Wold device for a certain family of probability measures: can we then deduce a similar device for mixtures of measures taken from that family? We give an affirmative answer to this question in certain cases, including the very important one of Gaussian mixtures.
Here, in more detail, is a road-map of the article.
Section 2 begins with a short introduction to the notion of the Cramér–Wold device. We then establish a rather general abstract result, Theorem 2.1, which shows how a Cramér–Wold theorem for a class of measures can be used to generate a similar theorem for mixtures drawn from the same class.
To apply Theorem 2.1, it is necessary to check that certain families of probability distributions are linearly independent, or equivalently, that they are identifiable. There is an extensive literature on techniques for proving identifiability. In Section 3, we review results concerning the identifiability of two particular families: the Gaussian distributions and the $t$-distributions. We also derive what we believe to be a new criterion for identifiability, based on a deep theorem in complex analysis due to Borel.
In Section 4, these ideas are applied to the important special case of Gaussian measures and Gaussian mixtures. In particular, we derive a Cramér–Wold theorem for Gaussian mixtures, Theorem 4.1, from the previously known one for Gaussian measures. It says that, if the one-dimensional projections of two mixtures of $\ell$ and $m$ normal distributions on $\mathbb{R}^{d}$ coincide on a certain pre-determined set of $(1/2)(l+m-1)(d^{2}+d-2)+1$ lines, then the two distributions are the same. A similar result, Theorem 4.2, is also established for mixtures of multivariate $t$-distributions.
We conclude in Section 5 with some remarks about potential applications of these results.
2. A Cramér–Wold theorem for mixtures
2.1. Introduction
Given a Borel probability measure $P$ on $\mathbb{R}^{d}$ and a vector subspace $H$ of $\mathbb{R}^{d}$, we write $P_{H}$ for the projection of $P$ onto $H$, namely the Borel probability measure on $H$ given by
$P_{H}(B):=P(\pi_{H}^{-1}(B)),$
where $\pi_{H}:\mathbb{R}^{d}\to H$ is the orthogonal projection of $\mathbb{R}^{d}$ onto $H$.
According to a well-known theorem of Cramér and Wold [3], if $P,Q$ are two Borel probability measures on $\mathbb{R}^{d}$, and if $P_{L}=Q_{L}$ for all lines $L$, then $P=Q$.
There have been numerous refinements of the Cramér–Wold theorem. To help describe these, it is convenient to introduce the following terminology. Let $\mathcal{P}$ be a family of Borel probability measures on $\mathbb{R}^{d}$ and let $\mathcal{H}$ be a family of vector subspaces of $\mathbb{R}^{d}$ (not necessarily all of the same dimension). We say that $\mathcal{H}$ is a Cramér–Wold system for $\mathcal{P}$ if, for every pair $P,Q\in\mathcal{P}$,
$P_{H}=Q_{H}~{}(\forall H\in\mathcal{H})\quad\Rightarrow\quad P=Q.$
In this terminology, the original Cramér–Wold theorem says simply that the family of all lines in $\mathbb{R}^{d}$ is a Cramér–Wold system for the family of Borel probability measures on $\mathbb{R}^{d}$. Here are some other examples.
  • If $\mathcal{P}$ is the family of compactly supported Borel probability measures on $\mathbb{R}^{2}$, then any infinite set of lines in $\mathbb{R}^{2}$ is a Cramér–Wold system for $\mathcal{P}$ (Rényi [17, Theorem 1]).
  • If $\mathcal{P}$ is the family of probability measures on $\mathbb{R}^{d}$ whose supports contain at most $k$ points, and if =$\mathcal{H}=\{H_{1},\dots,H_{k+1}\}$ is any family of $k+1$ subspaces of $\mathbb{R}^{d}$ such that $H_{i}^{\perp}\cap H_{j}^{\perp}=\{0\}$ whenever $i\neq j$, then $\mathcal{H}$ is a Cramér–Wold system for $\mathcal{P}$ (Heppes, [11, Theorem $1^{\prime}$]).
[…]
  • If $\mathcal{P}$ is the family of elliptical distributions on $\mathbb{R}^{d}$ and if $\mathcal{L}$ is the set of lines $\langle e_{i}+e_{j}\rangle~{}(1\leq i\leq j\leq d)$, where $\{e_{1},\dots,e_{d}\}$ is any basis of $\mathbb{R}^{d}$, then $\mathcal{L}$ is a Cramér–Wold system for $\mathcal{P}$ (Fraiman et al, [5, Theorem 1]).
Further results of this kind may be found in [1, 6, 7, 8, 10].
2.2. Mixtures
Let $d\geq 1$ and let $\mathcal{P}$ be a family of Borel probability measures on $\mathbb{R}^{d}$.
We say that $\mathcal{P}$ is linearly dependent if there exist distinct $P_{1},\dots,P_{n}\in\mathcal{P}$ and non-zero $\lambda_{1},\dots,\lambda_{n}\in\mathbb{R}$ such that $\sum_{j=1}^{n}\lambda_{j}P_{j}=0$. Otherwise $\mathcal{P}$ is linearly independent.
A $\mathcal{P}$-mixture is a convex combination of measures from $\mathcal{P}$, in other words, a measure of the form $\sum_{j=1}^{n}\lambda_{j}P_{j}$, where $P_{1},\dots,P_{n}\in\mathcal{P}$ and $\lambda_{1},\dots,\lambda_{n}\geq 0$ with $\sum_{j=1}^{n}\lambda_{j}=1$.
Our aim in this section is to establish the following Cramér–Wold theorem for $\mathcal{P}$-mixtures.
Theorem 2.1.
Let $\mathcal{P}$ be a family of Borel probability measures on $\mathbb{R}^{d}$, let $\mathcal{H}$ be a collection of vector subspaces of $\mathbb{R}^{d}$, and let $\ell,m\geq 1$. Suppose that:
  1. (i)
    For each $H\in\mathcal{H}$, the distinct measures in the set $\{P_{H}:P\in\mathcal{P}\}$ are linearly independent.
  2. (ii)
    Each partition of $\mathcal{H}$ into $\ell+m-1$ subsets contains at least one Cramér–Wold system for $\mathcal{P}$.
Let $P$ and $Q$ be convex combinations of $\ell$ and $m$ measures from $\mathcal{P}$ respectively. If $P_{H}=Q_{H}$ for all $H\in\mathcal{H}$, then $P=Q$.
Proof.
We argue by contradiction. Suppose that $P_{H}=Q_{H}$ for all $H\in\mathcal{H}$, but that $P\neq Q$. Then the difference $P-Q$ can be written as $P-Q=\sum_{j=1}^{k}\lambda_{j}P_{j}$, where $P_{1},\dots,P_{k}\in\mathcal{P}$ are distinct, where $\lambda_{1},\dots,\lambda_{k}\in\mathbb{R}$ are non-zero, and where $k\leq\ell+m$. Set
$\mathcal{H}_{j}:=\{H\in\mathcal{H}:(P_{1})_{H}=(P_{j})_{H}\}\quad(j=2,\dots,k).$
Clearly no $\mathcal{H}_{j}$ is Cramér–Wold system for $\mathcal{P}$. If we had $\cup_{j=2}^{k}\mathcal{H}_{j}=\mathcal{H}$, then we could construct a partition of $\mathcal{H}$ into $k-1$ sets (perhaps some of them empty) none of which is a Cramér–Wold system for $\mathcal{P}$, contradicting the assumption (ii) on $\mathcal{H}$. We conclude that $\cup_{j=2}^{k}\mathcal{H}_{j}\neq\mathcal{H}$, so there exists an $H_{0}\in\mathcal{H}$ such that $(P_{1})_{H_{0}}\neq(P_{j})_{H_{0}}~{}(j=2,\dots,k)$.
By assumption (i), the distinct measures in the set $\{(P_{j})_{H_{0}}:j=1,\dots d\}$ are linearly independent. In particular, since $(P_{1})_{H_{0}}\neq(P_{j})_{H_{0}}~{}(j=2,\dots,k)$, it follows that $(P_{1})_{H_{0}}$ is not in the span of $\{(P_{j})_{H_{0}}:j=2,\dots d\}$. On the other hand, we have
$\sum_{j=1}^{k}\lambda_{j}(P_{j})_{H_{0}}=\Bigl{(}\sum_{j=1}^{k}\lambda_{j}P_{j}\Bigr{)}_{H_{0}}=(P-Q)_{H_{0}}=P_{H_{0}}-Q_{H_{0}}=0.$
Since $\lambda_{1}\neq 0$, this is a contradiction. ∎
3. Linear independence and identifiability
3.1. Introduction
Condition (i) in Theorem 2.1 begs the question as to how one determines whether a set of measures on a subspace $H$ of $\mathbb{R}^{d}$ is linearly independent.
In the statistical literature, the notion of linear independence of measures is synonymous with that of identifiability. We say that a family $\mathcal{P}$ of Borel probability measures on $\mathbb{R}^{d}$ is identifiable if it is impossible to express any $\mathcal{P}$-mixture as two different convex combinations of elements of $\mathcal{P}$. It is easy to see that $\mathcal{P}$ is identifiable if and only if it is linearly independent (see e.g. [19, Theorem 3.1.1]). There is an extensive literature concerning techniques for proving identifiability/linear independence. A useful background reference is the book [19].
3.2. Linear independence of Gaussian and $t$-distributions
We shall need two results in particular. The first concerns the family of Gaussian distributions, with densities
(1) $f_{\mu,\sigma}(x)=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\Bigl{(}-\frac{(x-\mu)^{2}}{2\sigma^{2}}\Bigr{)}\quad(x\in\mathbb{R}).$
Theorem 3.1.
The family of Gaussian distributions $\{f_{\mu,\sigma}:\mu\in\mathbb{R},\sigma>0\}$ on $\mathbb{R}$ is linearly independent.
Proof.
This result is well known. See for example [19, Example 3.1.4]. ∎
The second result treats the identifiability/linear independence of the family of $t$-distributions. Recall that a $t$-distribution on $\mathbb{R}$ with $\nu$ degrees of freedom is a Borel measure with density of the form
$f_{\nu,\mu,\sigma}(x)=c_{\nu,\mu,\sigma}\Bigl{(}1+\frac{(x-\mu)^{2}}{\nu\sigma^{2}}\Bigr{)}^{-(\nu+1)/2},$
[…]
Theorem 3.2.
The family of $t$-distributions $\{f_{\nu,\mu,\sigma}:\nu\in\mathbb{Z}^{+},\mu\in\mathbb{R},\sigma>0\}$ is linearly independent.
Special cases of this result have previously been obtained by Otiniano et al. (case $\mu=0$), see [16, Proposition 4], and by Ho–Nguyen (case $\nu$ odd), see [12, Theorem 3.4(b)]. However, we have not been able to find the general case elsewhere in the literature, so we present a proof here.
Proof.
Suppose, on the contrary, that there is a linear relation of the form
(2) $\sum_{j=1}^{n}\lambda_{j}f_{\nu_{j},\mu_{j},\sigma_{j}}(x)=0\quad(x\in\mathbb{R}),$
where the triples $(\nu_{j},\mu_{j},\sigma_{j})$ are distinct and the coefficients $\lambda_{j}\neq 0$. Re-ordering the triples, we may further suppose that $\nu_{1}\geq\nu_{j}$ for all $j\geq 2$.
If all the $\nu_{j}$ are odd integers, then we may argue as follows. Each function $f_{\nu_{j},\mu_{j},\sigma_{j}}$ extends to a function in the complex plane $\mathbb{C}$
$f_{\nu_{j},\mu_{j},\sigma_{j}}(z)=c\Bigl{(}1+\frac{(z-\mu_{j})^{2}}{\nu_{j}\sigma_{j}^{2}}\Bigr{)}^{-(\nu_{j}+1)/2},$
which is holomorphic on $\mathbb{C}$ except for poles at the two points $\mu_{j}\pm i\sigma_{j}\sqrt{\nu_{j}}$. Moreover, the order of the poles at these points is exactly $(\nu_{j}+1)/2$. In particular, it follows that, if $j\geq 2$, then either $f_{\nu_{j},\mu_{j},\sigma_{j}}$ has no singularity at $\mu_{1}+i\sigma_{1}\sqrt{\nu_{1}}$, or it has a pole there of order strictly less than $\nu_{1}$. Thus
$\frac{f_{\nu_{j},\mu_{j},\sigma_{j}}(z)}{f_{\nu_{1},\mu_{1},\sigma_{1}}(z)}\to 0\quad\text{as~{}}z\to(\mu_{1}+i\sigma_{1}\sqrt{\nu_{1}})\quad(2\leq j\leq n).$
Now, by the identity principle, the relation (2) extends to
$\sum_{j=1}^{n}\lambda_{j}f_{\nu_{j},\mu_{j},\sigma_{j}}(z)=0\quad(z\in\mathbb{C},\,z\neq\mu_{j}\pm i\sigma_{j}\sqrt{\nu_{j}}).$
Dividing both sides by $f_{\nu_{1},\mu_{1},\sigma_{1}}(z)$ and then letting $z\to\mu_{1}+i\sigma_{1}\sqrt{\nu_{1}}$, we deduce that $\lambda_{1}=0$, contrary to hypothesis.
If some of the $\nu_{j}$ are even integers, then the argument needs to be adjusted slightly, since the presence of square roots means that the holomorphic extensions of $f_{\nu_{j},\mu_{j},\sigma_{j}}$ are multivalued. The problem is fixed as follows. We construct a finite set of non-intersecting half-lines terminating at the points $\mu_{j}\pm i\sigma_{j}\sqrt{\nu_{j}}$ and not crossing the real axis. The complement of these half-lines is then a simply connected domain $D$ containing $\mathbb{R}$ whose boundary contains each point $\mu_{j}\pm i\sigma_{j}\sqrt{\nu_{j}}$. Each function $f_{\nu_{j},\mu_{j},\sigma_{j}}$ has a single-valued holomorphic extension to $D$, and we can now argue exactly as before. ∎
3.3. Linear independence of measures via Borel’s theorem
Theorems 3.1 and 3.2 above are just two examples of a large class of theorems asserting the identifiability of various families of distributions. The proofs of many of these results are very similar, and indeed our proof of Theorem 3.2 relies on the same ideas. There is even an abstract formulation of the general method due to Teicher [18, Theorem 2].
In this subsection we present a further general result on identifiability/linear independence, based on a completely different principle, namely a classic result in complex analysis due to Borel [2]. We have not seen this theorem used elsewhere in the statistical literature, and we believe that the following application may be of interest.
Theorem 3.3.
Let $\mathcal{P}$ be a family of Borel probability measures on $\mathbb{R}^{d}$. Suppose that the characteristic function of every measure in $\mathcal{P}$ is the the restriction to $\mathbb{R}^{d}$ of a nowhere-vanishing holomorphic function on $\mathbb{C}^{d}$. Then $\mathcal{P}$ is linearly independent.
Remarks.
(i) This result immediately yields an alternative proof of Theorem 3.1, since the characteristic functions of the Gaussian distributions (1) have the form $\phi(\xi)=\exp(i\mu\xi-\sigma^{2}\xi^{2}/2)$.
[…]
(iii) Suppose that characteristic function of $P$ is the restriction to $\mathbb{R}^{d}$ of a holomorphic function $f$ on $\mathbb{C}^{d}$. A sufficient condition for $f$ to be zero-free on $\mathbb{C}^{d}$ is that $P$ be infinitely divisible. This follows from [13, Theorem 3.1] and [15, Theorem 4].
(iv) The following simple example shows that the nowhere-vanishing condition in Theorem 3.3 cannot be omitted. Let $P_{0}$ and $P_{1}$ be the Dirac measures on $\mathbb{R}$ concentrated at 00 and $1$ respectively, and let $P_{2}:=(1/3)P_{0}+(2/3)P_{1}$. The characteristic functions are the $P_{j}$ are given by $\phi_{P_{0}}(\xi)=1$ and $\phi_{P_{1}}(\xi)=e^{i\xi}$ and $\phi_{P_{2}}=(1/3)(1+2e^{i\xi})$. All three are nowhere-zero on $\mathbb{R}$ and all three extend to be holomorphic on $\mathbb{C}$. However, the set $\{P_{0},P_{1},P_{2}\}$ is clearly linearly dependent. Theorem 3.3 does not apply in this case, because the holomorphic extension of $\phi_{P_{2}}$ has zeros in $\mathbb{C}$ (namely at $i\log 2+2\pi n$ for each integer $n$).
As mentioned above, the proof of Theorem 3.3 is based on a theorem of Borel. The precise version that we need is due to Green [9, p.98]:
Lemma 3.4.
Let $g_{1},\dots,g_{n}:\mathbb{C}^{d}\to\mathbb{C}$ be holomorphic functions such that
$\exp(g_{1})+\cdots+\exp(g_{n})\equiv 0.$
Then, for some distinct $j,k$, the function $g_{j}-g_{k}$ is constant.
Proof of Theorem 3.3.
We argue by contradiction. Suppose that $\mathcal{P}$ is linearly dependent, so there exist distinct $P_{1},\dots,P_{n}\in\mathcal{P}$ and non-zero scalars $\lambda_{1},\dots,\lambda_{n}$ such that $\sum_{j=1}^{n}\lambda_{j}P_{j}=0$. Then the characteristic functions $\phi_{P_{j}}$ of the $P_{j}$ satisfy $\sum_{j=1}^{n}\lambda_{j}\phi_{P_{j}}(\xi)=0$ for all $\xi\in\mathbb{R}^{d}$. By assumption, each $\phi_{P_{j}}$ is the restriction to $\mathbb{R}^{d}$ of a nowhere-vanishing holomorphic function on $\mathbb{C}^{d}$. Thus we may write $\lambda_{j}\phi_{P_{j}}=\exp g_{j}|_{\mathbb{R}^{d}}$, where $g_{j}:\mathbb{C}^{d}\to\mathbb{C}$ is a holomorphic function on $\mathbb{C}^{d}$, and $\sum_{j=1}^{n}\exp g_{j}(\xi)=0$ for all $\xi\in\mathbb{R}^{d}$. By the identity principle, a holomorphic function on $\mathbb{C}^{d}$ that vanishes on $\mathbb{R}^{d}$ is identically zero on $\mathbb{C}^{d}$. Therefore $\sum_{j=1}^{n}\exp g_{j}(\zeta)=0$ for all $\zeta\in\mathbb{C}^{d}$. We now invoke Lemma 3.4, to deduce that $g_{j}-g_{k}$ is constant for some pair of distinct indices $j,k$. This implies that $\phi_{P_{j}}/\phi_{P_{k}}$ is constant on $\mathbb{R}^{d}$. As $P_{j},P_{k}$ are probability measures, we have $\phi_{P_{j}}(0)=\phi_{P_{k}}(0)=1$. Therefore $\phi_{P_{j}}=\phi_{P_{k}}$ on $\mathbb{R}^{d}$. By the uniqueness theorem for characteristic functions, it follows that $P_{j}=P_{k}$. This contradicts the fact that $P_{j}$ and $P_{k}$ are distinct measures. ∎
Remark.
Theorem 3.3 actually implies a stronger form of itself, as follows. Given $A\in M_{d}(\mathbb{R})$ (the set of $d\times d$ matrices) and $b\in\mathbb{R}^{d}$, let us write $P_{A,b}$ for the Borel probability measure on $\mathbb{R}^{d}$ defined by
(3) $P_{A,b}(B):=P\bigl{(}\{x\in\mathbb{R}^{d}:Ax+b\in B\}\bigr{)}.$
Suppose that the characteristic function of each measure in $\mathcal{P}$ is the restriction to $\mathbb{R}^{d}$ of a nowhere-vanishing function on $\mathbb{C}^{d}$. Then, not only is $\mathcal{P}$ linearly independent, but even the distinct measures in the family $\{P_{A,b}:P\in\mathcal{P},A\in M_{d}(\mathbb{R}),b\in\mathbb{R}^{d}\}$ are linearly independent. Indeed, a simple calculation shows that
$\phi_{P_{A,b}}(\xi)=e^{ib\cdot\xi}\phi_{P}(A^{T}\xi)\quad(\xi\in\mathbb{R}^{d}),$
so, if $\phi_{P}$ is the restriction to $\mathbb{R}^{d}$ of a nowhere-vanishing holomorphic function on $\mathbb{C}^{d}$, then the same is true of $\phi_{P_{A,b}}$. Theorem 3.3 now gives the result.
4. Gaussian mixtures
4.1. Introduction
A Gaussian measure $P$ on $\mathbb{R}^{d}$ is one whose density has the form
(4) $\frac{1}{(2\pi\det(\Sigma))^{d/2}}\exp\Bigl{(}-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)\Bigr{)}\quad(x\in\mathbb{R}^{d}),$
where $\mu\in\mathbb{\mathbb{R}}^{d}$, and where $\Sigma$ is a real $d\times d$ positive-definite matrix. A Gaussian mixture is a measure on $\mathbb{R}^{d}$ that is a finite convex combination of Gaussian measures.
Mixtures of multivariate Gaussian distributions have several nice properties. In particular, in Titterington et al. [19], it is shown that Gaussian kernel density estimators can approximate any continuous density given enough kernels (universal consistency). It is well known that Gaussian mixtures are weak*-dense in the space of all Borel probability measures on $\mathbb{R}^{d}$. They also have numerous applications in statistics; for more on this, see Section 5 below.
4.2. A Cramér–Wold theorem for Gaussian mixtures
In this section we consider the problem of testing for equality for two Gaussian mixtures by looking at a finite number of projections. The basic theorem underlying this approach has two ingredients. One is the abstract result Theorem 2.1. The other is a characterization of Cramér–Wold systems for Gaussian measures in $\mathbb{R}^{d}$ (and, more generally, for elliptical distributions) established in [5, Theorems 1 and 2], which we now recall.
Let $S$ be a set of vectors in $\mathbb{R}^{d}$. Then the corresponding set of lines $\{\langle x\rangle:x\in S\}$ is a Cramér–Wold system for the Gaussian measures in $\mathbb{R}^{d}$ if and only if $S$ has the property that the only real symmetric $d\times d$ matrix $A$ satisfying $x^{T}Ax=0$ for all $x\in S$ is the zero matrix. A set $S$ with this property is called a symmetric-matrix uniqueness set (or sm-uniqueness set for short).
It was shown in [5] that an sm-uniqueness set for $\mathbb{R}^{d}$ spans $\mathbb{R}^{d}$ and that it contains at least $(d^{2}+d)/2$ vectors. We shall call $S$ a strong sm-uniqueness set if every subset of $S$ containing $(d^{2}+d)/2$ vectors is an sm-uniqueness set. We can now state our Cramér–Wold theorem for Gaussian mixtures.
Theorem 4.1.
Let $P$ and $Q$ be convex combinations of $\ell$ and $m$ Gaussian measures on $\mathbb{R}^{d}$ respectively. Let $S$ be a strong sm-uniqueness set for $\mathbb{R}^{d}$ containing at least $(1/2)(\ell+m-1)(d^{2}+d-2)+1$ vectors. If $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, then $P=Q$.
Proof.
We apply Theorem 2.1 with $\mathcal{P}$ equal to the set of Gaussian measures on $\mathbb{R}^{d}$ and :=$\mathcal{H}:=\{\langle x\rangle:x\in S\}$. All we need to do is to check that $\mathcal{P}$ and $\mathcal{H}$ satisfy the hypotheses (i) and (ii) in Theorem 2.1.
Concerning hypothesis (i), the projection of a multivariate Gaussian measure onto a line is just a Gaussian measure on that line. We have already seen in Theorem 3.1 that the set of all one-dimensional Gaussian measures is a linearly independent family. So hypothesis (i) holds.
As for hypothesis (ii), we argue as follows. If $S$ is partitioned into $\ell+m-1$ sets, then one of them, $S_{0}$ say, must contain at least $(d^{2}+d)/2$ vectors (otherwise $S$ would contain at most $(1/2)(\ell+m-1)(d^{2}+d-2)$ vectors, contrary to assumption). As $S$ is a strong sm-uniqueness set, it follows that $S_{0}$ is an sm-uniqueness set. In summary, if $S$ is partitioned into $\ell+m-1$ sets, then at least one of them is an sm-uniqueness set. In other words, if $\mathcal{H}$ is partitioned into $\ell+m-1$ sets, then at least one of them is a Cramér–Wold system for the family of Gaussian measures. Thus hypothesis (ii) holds, and we are done. ∎
4.3. A Cramér–Wold theorem for $t$-mixtures
We now establish an analogue of Theorem 4.1 for mixtures of multivariate $t$-distributions, thereby allowing heavy-tailed distributions. A $t$-distribution on $\mathbb{R}^{d}$ is a measure with density of the form
$f_{\nu,\mu,\Sigma}(x)=c_{\nu,\mu,\Sigma}\Bigl{(}1+\frac{(x-\mu)^{T}\Sigma^{-1}(x-\mu)}{\nu}\Bigr{)}^{-(\nu+d)/2}\quad(x\in\mathbb{R}^{d}),$
[…]
Theorem 4.2.
Let $P$ and $Q$ be convex combinations of $\ell$ and $m$ multivariate $t$-distributions on $\mathbb{R}^{d}$ respectively. Let $S$ be a strong sm-uniqueness set for $\mathbb{R}^{d}$ containing at least $(1/2)(\ell+m-1)(d^{2}+d-2)+1$ vectors. If $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, then $P=Q$.
Proof.
This is virtually identical to the proof of Theorem 4.1. Once again, we apply Theorem 2.1, and we need to check that the hypotheses (i) and (ii) of that theorem hold.
Hypothesis (i) holds because the one-dimensional projection of a multivariate $t$-distribution is a univariate $t$-distribution, and by Theorem 3.2 the family of all univariate $t$-distributions is a linearly independent set.
Hypothesis (ii) holds for the same reason that it did before. Indeed, [5, Theorem 1] applies to all elliptical distributions, which includes $t$-distributions as well as Gaussian ones. ∎
4.4. Strong sm-uniqueness sets
Theorems 4.1 and 4.2 beg the question as to whether there exist strong sm-uniqueness sets of arbitrarily large cardinality. The following result provides an affirmative answer, and suggests a realistic method for generating them.
Theorem 4.3.
Let $d\geq 2$, let $k\geq(d^{2}+d)/2$, and let
[…]
Then $V$ is an open subset of $\mathbb{R}^{dk}$, and $\mathbb{R}^{dk}\setminus V$ has Lebesgue measure zero.
Thus, if $v_{1},\dots,v_{k}$ are independent random vectors in $\mathbb{R}^{d}$ with distributions given by densities on $\mathbb{R}^{d}$, then, with probability one, the set $\{v_{1},\dots,v_{k}\}$ is a strong sm-uniqueness set for $\mathbb{R}^{d}$. To test whether a specific family $\{v_{1},\dots,v_{k}\}$ is a strong sm-uniqueness set, one can use the following criterion for sm-uniqueness sets, which is also an ingredient in the proof of Theorem 4.3.
Let $d\geq 2$ and set $D:=(d^{2}+d)/2$. Given $x=(t_{1},\dots,t_{d})\in\mathbb{R}^{d}$, let $\widehat{x}$ be the upper triangular $d\times d$ matrix with entries $\widehat{x}_{ij}:=t_{i}t_{j}~{}(1\leq i\leq j\leq d)$, but viewed as a column vector in $\mathbb{R}^{D}$.
Lemma 4.4.
A $D$-tuple $(x_{1},\dots,x_{D})$ of vectors in $\mathbb{R}^{d}$ is an sm-uniqueness set if and only if $\widehat{x}_{1},\dots,\widehat{x}_{D}$ are linearly independent vectors in $\mathbb{R}^{D}$, in other words, if and only if the determinant of the $D\times D$ block matrix $(\widehat{x}_{1}|\widehat{x}_{2}|\cdots|\widehat{x}_{D})$ does not vanish.
Proof.
This is [5, Corollary 5]. ∎
Proof of Theorem 4.3.
By Lemma 4.4, the set $\mathbb{R}^{dk}\setminus V$ can be expressed as the union
$\bigcup_{1\leq j_{1}<j_{2}<\dots<j_{D}\leq k}\Bigl{\{}((v_{1},\dots,v_{k})\in(\mathbb{R}^{d})^{k}:\det\bigl{(}\widehat{v}_{j_{1}}|\widehat{v}_{j_{2}}|\cdots|\widehat{v}_{j_{D}}\bigr{)}=0\Bigr{\}}.$
For each choice of $(j_{1},j_{2},\dots,j_{D})$, the map $(v_{1},\dots,v_{k})\mapsto\det(\widehat{v}_{j_{1}}|\widehat{v}_{j_{2}}|\cdots|\widehat{v}_{j_{D}})$ is a polynomial in the entries of $(v_{1},\dots,v_{k})$ that is not identically zero, so its zero set is a closed subset of $\mathbb{R}^{dk}$ of Lebesgue measure zero. As $\mathbb{R}^{dk}\setminus V$ is a finite union of such sets, it too is a closed subset of $\mathbb{R}^{dk}$ of Lebesgue measure zero. ∎
5. Concluding remarks
Gaussian-mixture models have been shown to be very effective in modeling different real data, see for instance Titterington et al. [19] for a deep study of their properties. They provide flexible and general models, and relevant applications can be found in the literature in different fields like density estimation, machine learning and clustering, among others. The estimation of these models is quite involved, in particular for high-dimensional data, typically using Markov-chain Monte-Carlo methods in a Bayesian framework. We believe that the Cramér–Wold device may well have a role to play in this circle of ideas, in particular through Theorems 3.1 and 3.2. We intend to explore this in a future paper.

### v2

@@
 In Section 3, we review results concerning the identifiability of two particular families: the Gaussian distributions and the $t$-distributions.
-We also derive what we believe to be a new criterion for identifiability, based on a deep theorem in complex analysis due to Borel.
+We also derive what we believe to be a new criterion for identifiability, based on a generalization of Picard’s theorem from complex analysis due to Borel.
 In Section 4, these ideas are applied to the important special case of Gaussian measures and Gaussian mixtures.
 In particular, we derive a Cramér–Wold theorem for Gaussian mixtures, Theorem 4.1, from the previously known one for Gaussian measures.
-It says that, if the one-dimensional projections of two mixtures of $\ell$ and $m$ normal distributions on $\mathbb{R}^{d}$ coincide on a certain pre-determined set of $(1/2)(l+m-1)(d^{2}+d-2)+1$ lines, then the two distributions are the same.
+It says that, if $P$ and $Q$ are both mixtures of $m$ normal distributions on $\mathbb{R}^{d}$, and if their one-dimensional projections coincide on a certain pre-determined set of $(2m-1)(d^{2}+d-2)/2+1$ lines, then $P$ and $Q$ are the same.
 A similar result, Theorem 4.2, is also established for mixtures of multivariate $t$-distributions.
-We conclude in Section 5 with some remarks about potential applications of these results.
+We conclude in Section 5 with some remarks about applications of these results.
 2. A Cramér–Wold theorem for mixtures
@@
 $P_{H}(B):=P(\pi_{H}^{-1}(B)),$
-where $\pi_{H}:\mathbb{R}^{d}\to H$ is the orthogonal projection of $\mathbb{R}^{d}$ onto $H$.
+where $B$ is an arbitrary Borel subset of $H$ and $\pi_{H}:\mathbb{R}^{d}\to H$ is the orthogonal projection of $\mathbb{R}^{d}$ onto $H$.
 According to a well-known theorem of Cramér and Wold [3], if $P,Q$ are two Borel probability measures on $\mathbb{R}^{d}$, and if $P_{L}=Q_{L}$ for all lines $L$, then $P=Q$.
@@
   • If $\mathcal{P}$ is the family of compactly supported Borel probability measures on $\mathbb{R}^{2}$, then any infinite set of lines in $\mathbb{R}^{2}$ is a Cramér–Wold system for $\mathcal{P}$ (Rényi [17, Theorem 1]).
-  • If $\mathcal{P}$ is the family of probability measures on $\mathbb{R}^{d}$ whose supports contain at most $k$ points, and if =$\mathcal{H}=\{H_{1},\dots,H_{k+1}\}$ is any family of $k+1$ subspaces of $\mathbb{R}^{d}$ such that $H_{i}^{\perp}\cap H_{j}^{\perp}=\{0\}$ whenever $i\neq j$, then $\mathcal{H}$ is a Cramér–Wold system for $\mathcal{P}$ (Heppes, [11, Theorem $1^{\prime}$]).
+  • If $\mathcal{P}$ is the family of probability measures on $\mathbb{R}^{d}$ whose supports contain at most $k$ points, and if =$\mathcal{H}=\{H_{1},\dots,H_{k+1}\}$ is any family of $k+1$ subspaces of $\mathbb{R}^{d}$ such that $H_{i}^{\perp}\cap H_{j}^{\perp}=\{0\}$ whenever $i\neq j$, then $\mathcal{H}$ is a Cramér–Wold system for $\mathcal{P}$ (Heppes, [12, Theorem $1^{\prime}$]).
 […]
@@
 Theorem 2.1.
-Let $\mathcal{P}$ be a family of Borel probability measures on $\mathbb{R}^{d}$, let $\mathcal{H}$ be a collection of vector subspaces of $\mathbb{R}^{d}$, and let $\ell,m\geq 1$. Suppose that:
+Let $\mathcal{P}$ be a family of Borel probability measures on $\mathbb{R}^{d}$, let $\mathcal{H}$ be a collection of vector subspaces of $\mathbb{R}^{d}$, and let $m\geq 1$. Suppose that:
   1. (i)
@@
   2. (ii)
-    Each partition of $\mathcal{H}$ into $\ell+m-1$ subsets contains at least one Cramér–Wold system for $\mathcal{P}$.
-Let $P$ and $Q$ be convex combinations of $\ell$ and $m$ measures from $\mathcal{P}$ respectively. If $P_{H}=Q_{H}$ for all $H\in\mathcal{H}$, then $P=Q$.
-Proof.
+    Each partition of $\mathcal{H}$ into $12m-$ 1 subsets contains at least one Cramér–Wold system for $\mathcal{P}$.
+Let $P$ and $Q$ each be convex combinations of $m$ measures from $\mathcal{P}$. If $P_{H}=Q_{H}$ for all $H\in\mathcal{H}$, then $P=Q$.
+Remark.
+The statement of Theorem 2.1 immediately raises the question of how to verify the hypotheses (i) and (ii).
+Hypothesis (i) will be treated in detail Section 3.
+Hypothesis (ii) will be addressed (at least in the special case of Gaussian and $t$-distributions) in Section 4.
+Proof of Theorem 2.1.
 We argue by contradiction.
 Suppose that $P_{H}=Q_{H}$ for all $H\in\mathcal{H}$, but that $P\neq Q$.
-Then the difference $P-Q$ can be written as $P-Q=\sum_{j=1}^{k}\lambda_{j}P_{j}$, where $P_{1},\dots,P_{k}\in\mathcal{P}$ are distinct, where $\lambda_{1},\dots,\lambda_{k}\in\mathbb{R}$ are non-zero, and where $k\leq\ell+m$.
+Then the difference $P-Q$ can be written as $P-Q=\sum_{j=1}^{k}\lambda_{j}P_{j}$, where $P_{1},\dots,P_{k}\in\mathcal{P}$ are distinct, where $\lambda_{1},\dots,\lambda_{k}\in\mathbb{R}$ are non-zero, and where $k\leq 2m$.
 Set
 $\mathcal{H}_{j}:=\{H\in\mathcal{H}:(P_{1})_{H}=(P_{j})_{H}\}\quad(j=2,\dots,k).$
-Clearly no $\mathcal{H}_{j}$ is Cramér–Wold system for $\mathcal{P}$.
+Clearly no $\mathcal{H}_{j}$ is a Cramér–Wold system for $\mathcal{P}$.
 If we had $\cup_{j=2}^{k}\mathcal{H}_{j}=\mathcal{H}$, then we could construct a partition of $\mathcal{H}$ into $k-1$ sets (perhaps some of them empty) none of which is a Cramér–Wold system for $\mathcal{P}$, contradicting the assumption (ii) on $\mathcal{H}$.
 We conclude that $\cup_{j=2}^{k}\mathcal{H}_{j}\neq\mathcal{H}$, so there exists an $H_{0}\in\mathcal{H}$ such that $(P_{1})_{H_{0}}\neq(P_{j})_{H_{0}}~{}(j=2,\dots,k)$.
-By assumption (i), the distinct measures in the set $\{(P_{j})_{H_{0}}:j=1,\dots d\}$ are linearly independent.
+By assumption (i), the distinct measures in the set $\{(P_{j})_{H_{0}}:j=1,\dots k\}$ are linearly independent.
 In particular, since $(P_{1})_{H_{0}}\neq(P_{j})_{H_{0}}~{}(j=2,\dots,k)$, it follows that $(P_{1})_{H_{0}}$ is not in the span of $\{(P_{j})_{H_{0}}:j=2,\dots d\}$.
@@
 Proof.
-This result is well known. See for example [19, Example 3.1.4]. ∎
+This result is well known. An early reference is [20, Proposition 2]. See also [19, Example 3.1.4]. ∎
 The second result treats the identifiability/linear independence of the family of $t$-distributions.
@@
 The family of $t$-distributions $\{f_{\nu,\mu,\sigma}:\nu\in\mathbb{Z}^{+},\mu\in\mathbb{R},\sigma>0\}$ is linearly independent.
-Special cases of this result have previously been obtained by Otiniano et al.
-(case $\mu=0$), see [16, Proposition 4], and by Ho–Nguyen (case $\nu$ odd), see [12, Theorem 3.4(b)].
-However, we have not been able to find the general case elsewhere in the literature, so we present a proof here.
 Proof.
-Suppose, on the contrary, that there is a linear relation of the form
-(2) $\sum_{j=1}^{n}\lambda_{j}f_{\nu_{j},\mu_{j},\sigma_{j}}(x)=0\quad(x\in\mathbb{R}),$
-where the triples $(\nu_{j},\mu_{j},\sigma_{j})$ are distinct and the coefficients $\lambda_{j}\neq 0$. Re-ordering the triples, we may further suppose that $\nu_{1}\geq\nu_{j}$ for all $j\geq 2$.
-If all the $\nu_{j}$ are odd integers, then we may argue as follows. Each function $f_{\nu_{j},\mu_{j},\sigma_{j}}$ extends to a function in the complex plane $\mathbb{C}$
-$f_{\nu_{j},\mu_{j},\sigma_{j}}(z)=c\Bigl{(}1+\frac{(z-\mu_{j})^{2}}{\nu_{j}\sigma_{j}^{2}}\Bigr{)}^{-(\nu_{j}+1)/2},$
-which is holomorphic on $\mathbb{C}$ except for poles at the two points $\mu_{j}\pm i\sigma_{j}\sqrt{\nu_{j}}$.
-Moreover, the order of the poles at these points is exactly $(\nu_{j}+1)/2$.
-In particular, it follows that, if $j\geq 2$, then either $f_{\nu_{j},\mu_{j},\sigma_{j}}$ has no singularity at $\mu_{1}+i\sigma_{1}\sqrt{\nu_{1}}$, or it has a pole there of order strictly less than $\nu_{1}$.
-Thus
-$\frac{f_{\nu_{j},\mu_{j},\sigma_{j}}(z)}{f_{\nu_{1},\mu_{1},\sigma_{1}}(z)}\to 0\quad\text{as~{}}z\to(\mu_{1}+i\sigma_{1}\sqrt{\nu_{1}})\quad(2\leq j\leq n).$
-Now, by the identity principle, the relation (2) extends to
-$\sum_{j=1}^{n}\lambda_{j}f_{\nu_{j},\mu_{j},\sigma_{j}}(z)=0\quad(z\in\mathbb{C},\,z\neq\mu_{j}\pm i\sigma_{j}\sqrt{\nu_{j}}).$
-Dividing both sides by $f_{\nu_{1},\mu_{1},\sigma_{1}}(z)$ and then letting $z\to\mu_{1}+i\sigma_{1}\sqrt{\nu_{1}}$, we deduce that $\lambda_{1}=0$, contrary to hypothesis.
-If some of the $\nu_{j}$ are even integers, then the argument needs to be adjusted slightly, since the presence of square roots means that the holomorphic extensions of $f_{\nu_{j},\mu_{j},\sigma_{j}}$ are multivalued.
-The problem is fixed as follows.
-We construct a finite set of non-intersecting half-lines terminating at the points $\mu_{j}\pm i\sigma_{j}\sqrt{\nu_{j}}$ and not crossing the real axis.
-The complement of these half-lines is then a simply connected domain $D$ containing $\mathbb{R}$ whose boundary contains each point $\mu_{j}\pm i\sigma_{j}\sqrt{\nu_{j}}$.
-Each function $f_{\nu_{j},\mu_{j},\sigma_{j}}$ has a single-valued holomorphic extension to $D$, and we can now argue exactly as before.
-∎
+See for example [13, Section 3, Example 1]. ∎
 3.3. Linear independence of measures via Borel’s theorem
 Theorems 3.1 and 3.2 above are just two examples of a large class of theorems asserting the identifiability of various families of distributions.
-The proofs of many of these results are very similar, and indeed our proof of Theorem 3.2 relies on the same ideas.
-There is even an abstract formulation of the general method due to Teicher [18, Theorem 2].
-In this subsection we present a further general result on identifiability/linear independence, based on a completely different principle, namely a classic result in complex analysis due to Borel [2].
+The proofs of many of these results are very similar, and there is even an abstract formulation of the general method due to Teicher [18, Theorem 2].
+In this subsection we present a further general result on identifiability/linear independence, based on a completely different principle, namely a generalization of Picard’s theorem from complex analysis due to Borel [2].
 We have not seen this theorem used elsewhere in the statistical literature, and we believe that the following application may be of interest.
@@
 Let $\mathcal{P}$ be a family of Borel probability measures on $\mathbb{R}^{d}$.
-Suppose that the characteristic function of every measure in $\mathcal{P}$ is the the restriction to $\mathbb{R}^{d}$ of a nowhere-vanishing holomorphic function on $\mathbb{C}^{d}$.
+Suppose that the characteristic function of every measure in $\mathcal{P}$ is the restriction to $\mathbb{R}^{d}$ of a nowhere-vanishing holomorphic function on $\mathbb{C}^{d}$.
 Then $\mathcal{P}$ is linearly independent.
@@
 A sufficient condition for $f$ to be zero-free on $\mathbb{C}^{d}$ is that $P$ be infinitely divisible.
-This follows from [13, Theorem 3.1] and [15, Theorem 4].
+This follows from [14, Theorem 3.1] and [16, Theorem 4].
 (iv) The following simple example shows that the nowhere-vanishing condition in Theorem 3.3 cannot be omitted.
@@
 However, the set $\{P_{0},P_{1},P_{2}\}$ is clearly linearly dependent.
-Theorem 3.3 does not apply in this case, because the holomorphic extension of $\phi_{P_{2}}$ has zeros in $\mathbb{C}$ (namely at $i\log 2+2\pi n$ for each integer $n$).
-As mentioned above, the proof of Theorem 3.3 is based on a theorem of Borel. The precise version that we need is due to Green [9, p.98]:
+Theorem 3.3 does not apply in this case, because the holomorphic extension of $\phi_{P_{2}}$ has zeros in $\mathbb{C}$ (namely at $i\log 2+(2n+1)\pi$ for each integer $n$).
+As mentioned above, the proof of Theorem 3.3 is based on a theorem of Borel. The precise version that we need is due to Green [10, p.98]:
 Lemma 3.4.
@@
 A Gaussian measure $P$ on $\mathbb{R}^{d}$ is one whose density has the form
-(4) $\frac{1}{(2\pi\det(\Sigma))^{d/2}}\exp\Bigl{(}-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)\Bigr{)}\quad(x\in\mathbb{R}^{d}),$
+(3) $\frac{1}{\det(2\pi\Sigma)^{1/2}}\exp\Bigl{(}-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)\Bigr{)}\quad(x\in\mathbb{R}^{d}),$
 where $\mu\in\mathbb{\mathbb{R}}^{d}$, and where $\Sigma$ is a real $d\times d$ positive-definite matrix.
@@
 We shall call $S$ a strong sm-uniqueness set if every subset of $S$ containing $(d^{2}+d)/2$ vectors is an sm-uniqueness set.
+A method for generating examples of strong sm-uniqueness sets will be described in Section 4.4 below.
 We can now state our Cramér–Wold theorem for Gaussian mixtures.
 Theorem 4.1.
-Let $P$ and $Q$ be convex combinations of $\ell$ and $m$ Gaussian measures on $\mathbb{R}^{d}$ respectively.
-Let $S$ be a strong sm-uniqueness set for $\mathbb{R}^{d}$ containing at least $(1/2)(\ell+m-1)(d^{2}+d-2)+1$ vectors.
+Let $P$ and $Q$ each be convex combinations of $m$ Gaussian measures on $\mathbb{R}^{d}$.
+Let $S$ be a strong sm-uniqueness set for $\mathbb{R}^{d}$ containing at least $(2m-1)(d^{2}+d-2)/2+1$ vectors.
 If $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, then $P=Q$.
@@
 As for hypothesis (ii), we argue as follows.
-If $S$ is partitioned into $\ell+m-1$ sets, then one of them, $S_{0}$ say, must contain at least $(d^{2}+d)/2$ vectors (otherwise $S$ would contain at most $(1/2)(\ell+m-1)(d^{2}+d-2)$ vectors, contrary to assumption).
+If $S$ is partitioned into $12m-$ 1 sets, then one of them, $S_{0}$ say, must contain at least $(d^{2}+d)/2$ vectors (otherwise $S$ would contain at most $(2m-1)(d^{2}+d-2)/2$ vectors, contrary to assumption).
 As $S$ is a strong sm-uniqueness set, it follows that $S_{0}$ is an sm-uniqueness set.
-In summary, if $S$ is partitioned into $\ell+m-1$ sets, then at least one of them is an sm-uniqueness set.
-In other words, if $\mathcal{H}$ is partitioned into $\ell+m-1$ sets, then at least one of them is a Cramér–Wold system for the family of Gaussian measures.
+In summary, if $S$ is partitioned into $12m-$ 1 sets, then at least one of them is an sm-uniqueness set.
+In other words, if $\mathcal{H}$ is partitioned into $12m-$ 1 sets, then at least one of them is a Cramér–Wold system for the family of Gaussian measures.
 Thus hypothesis (ii) holds, and we are done.
@@
 Theorem 4.2.
-Let $P$ and $Q$ be convex combinations of $\ell$ and $m$ multivariate $t$-distributions on $\mathbb{R}^{d}$ respectively.
-Let $S$ be a strong sm-uniqueness set for $\mathbb{R}^{d}$ containing at least $(1/2)(\ell+m-1)(d^{2}+d-2)+1$ vectors.
+Let $P,Q$ each be convex combinations of $m$ multivariate $t$-distributions on $\mathbb{R}^{d}$.
+Let $S$ be a strong sm-uniqueness set for $\mathbb{R}^{d}$ containing at least $(2m-1)(d^{2}+d-2)/2+1$ vectors.
 If $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, then $P=Q$.
@@
 ∎
+4.5. Two examples
+In this subsection, we examine the rationale for the number $(2m-1)(d^{2}+d-2)/2+1$ appearing in Theorems 4.1 and 4.2.
+In particular, we explain the quadratic growth in $d$ and the linear growth in $m$.
+The quadratic growth in $d$ is easily justified by the following theorem, essentially taken from [5].
+We consider the special case $m=1$ (so that $P,Q$ are no longer mixtures).
+Thus $(2m-1)(d^{2}+d-2)/2+1=(d^{2}+d)/2$.
+Theorem 4.5.
+Let $d\geq 2$ and let $S$ be any set of vectors in $\mathbb{R}^{d}$ containing strictly fewer than $(d^{2}+d)/2$ vectors.
+Then there exist Gaussian measures $P,Q$ on $\mathbb{R}^{d}$ such that $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, but $P\neq Q$.
+Proof.
+By [5, Proposition 2], the fact that $S$ has fewer than $(d^{2}+d)/2$ elements implies that it is not an sm-uniqueness set. The existence of $P$ and $Q$ now follows from [5, Theorem 2]. ∎
+This result explains not only the presence of the factor $(d^{2}+d)/2$ in Theorem 4.1, but also the need for considering of the notion of sm-uniqueness set.
+We remark in passing that the quantity $(d^{2}+d)/2$ arises here because it is the number of degrees of freedom of a $d\times d$ covariance matrix.
+This is discussed in more detail in [5, Section 2].
+Now we turn to the dependence on $m$, namely the number of Gaussian (or $t$-) distributions making up the mixtures $P$ and $Q$.
+In our first example, we take $d=2$ and $m=2$.
+Though apparently modest, we believe that this example is instructive.
+Example 4.6.
+Let $S$ be the subset of $\mathbb{R}^{2}$ defined by
+$S:=\Bigl{\{}{1\choose 0},{0\choose 1},{1\choose 1},{1\choose-1}\Bigr{\}}.$
+Then $S$ is a strong sm-uniqueness set, and there exist measures $P,Q$ on $\mathbb{R}^{2}$, each a convex combination of two Gaussian measures, such that $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, but $P\neq Q$.
+Proof.
+First of all, we remark that, by [5, Corollary 3], a set of three vectors in $\mathbb{R}^{2}$ is an sm-uniqueness set provided that no two vectors in the set are multiples of one another.
+It follows that $S$ is a strong sm-uniqueness set.
+Now set
+$\Sigma_{1}:=\begin{pmatrix}2&0\\ 0&1\end{pmatrix},\quad\Sigma_{2}:=\begin{pmatrix}1&1\\ 1&2\end{pmatrix},\quad\Sigma_{1}^{\prime}:=\begin{pmatrix}2&1\\ 1&1\end{pmatrix},\quad\Sigma_{2}^{\prime}:=\begin{pmatrix}1&0\\ 0&2\end{pmatrix}.$
+Let $P:=(P_{1}+P_{2})/2$, where $P_{j}$ is Gaussian with mean 00 and covariance $\Sigma_{j}$, and let $Q:=(Q_{1}+Q_{2})/2$, where $Q_{j}$ is Gaussian with mean 00 and covariance $\Sigma_{j}^{\prime}$.
+Clearly $P\neq Q$.
+Elementary calculations show that, if $x={1\choose 0}$ or (${0\choose 1}$, then
+$x^{T}\Sigma_{1}x=x^{T}\Sigma_{1}^{\prime}x\quad\text{and}\quad x^{T}\Sigma_{2}x=x^{T}\Sigma_{2}^{\prime}x,$
+while if $x={1\choose 1}$ or (${1\choose-1}$, then
+$x^{T}\Sigma_{1}x=x^{T}\Sigma_{2}^{\prime}x\quad\text{and}\quad x^{T}\Sigma_{2}x=x^{T}\Sigma_{1}^{\prime}x.$
+Note that, in this last display, $\Sigma_{1}^{\prime}$ and $\Sigma_{2}^{\prime}$ have changed places.
+However, since $P$ and $Q$ are averages of the $P_{j}$ and $Q_{j}$ respectively, we do have $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$.
+∎
+If $d=2$ and $m=2$, then $(2m-1)(d^{2}+d-2)/2+1=7$.
+Example 4.6 shows that, in Theorem 4.1, this number cannot be replaced by $4$.
+However, calculations similar to those carried out in Example 4.6 indicate that $5$ lines do suffice.
+Thus, in this case at least, the bound $(2m-1)(d^{2}+d-2)/2+1$ is not sharp.
+Despite the lack of sharpness, the linear growth in $m$ is correct.
+To see this, we consider a second example where, instead of varying the covariance matrices, we vary the means.
+This example is inspired by one for discrete measures presented in [12].
+Example 4.7.
+Let $m\geq 2$ and define a subset $S$ of $\mathbb{R}^{2}$ by
+$S:=\Bigl{\{}\begin{pmatrix}\cos(j\pi/m)\\ \sin(j\pi/m)\end{pmatrix}:j=1,\dots,m\Bigr{\}}.$
+Then $S$ is a strong sm-uniqueness set, and there exist measures $P,Q$ on $\mathbb{R}^{2}$, each a convex combination of $m$ Gaussian measures, such that $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, but $P\neq Q$.
+Proof.
+As in the previous example, a set of three vectors in $\mathbb{R}^{2}$ is an sm-uniqueness set provided that no two vectors in the set are multiples of one another.
+It follows that $S$ is a strong sm-uniqueness set.
+For $j=1,\dots,m$, set
+$\mu_{j}:=\begin{pmatrix}\cos((2j-1/2)\pi/m)\\ \sin((2j-1/2)\pi/m)\end{pmatrix}\quad\text{and}\quad\mu_{j}^{\prime}:=\begin{pmatrix}\cos((2j+1/2)\pi/m)\\ \sin((2j+1/2)\pi/m)\end{pmatrix}.$
+Thus the sequence $\mu_{1},\mu_{1}^{\prime},\mu_{2},\mu_{2}^{\prime},\dots,\mu_{m},\mu_{m}^{\prime}$ traces out, in order, the vertices of a regular $(2m)$-gon centred at the origin.
+The lines defined by the set $S$ are precisely the bisectors of this $(2m)$-gon passing through the midpoints of the edges.
+For $j=1,\dots,m$, let $P_{j},Q_{j}$ be Gaussian measures with means at $\mu_{j},\mu_{j}^{\prime}$ respectively, all of them having covariance matrix equal to the identity.
+Finally, let $P=(1/m)\sum_{j=1}^{m}P_{j}$ and $Q:=(1/m)\sum_{j=1}^{m}Q_{j}$.
+Then it is geometrically clear that $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, but evidently $P\neq Q$.
+∎
+This example works because of the symmetry of the set-up.
+Though such symmetry is rather exceptional, there is nothing in Theorem 4.1 to exclude this possibility, and so it needs to be taken into account when deciding the number of lines to be used.
+Example 4.7 shows that (when $d=2$), this is number has to be at least $m+1$.
+This explains the linear growth in $m$ in the bound $(2m-1)(d^{2}+d-2)/2+1$.
 5. Concluding remarks
@@
 The estimation of these models is quite involved, in particular for high-dimensional data, typically using Markov-chain Monte-Carlo methods in a Bayesian framework.
-We believe that the Cramér–Wold device may well have a role to play in this circle of ideas, in particular through Theorems 3.1 and 3.2.
-We intend to explore this in a future paper.
+The Cramér–Wold device has a role to play in this circle of ideas, in particular through Theorems 4.1 and 4.2.
+This is explored in detail in the companion paper [8].